OpenStudy (anonymous):

what is the average (arithmetic mean) of x^2+2x-3, 3x^2-2x-3 ,and 30-4x^2?

OpenStudy (kropot72):

The sum of the three expressions is 24. The terms in x cancel in the addition. Therefore the AM is 24/3 = 8

OpenStudy (bradely):

Arithmetic mean= ((x^2+2x-3)+( 3x^2-2x-3)+(30-4x^2))/3 =24/3 =8